OpenStudy (anonymous):

How do you find the integral of x^3*sinx^2 dx?

OpenStudy (anonymous):

u= x^2 du = 2xdx so (1/2)du= xdx. and also break up your integral as x*x^2sin(x^2)dx so your substitution would look like (1/2)usin(u)du and now apply integration by parts

OpenStudy (anonymous):

Thank you so much! :)